Session: Sep 1999 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Cattani at Mon Sep 20 20:15:00 METDST 1999) Minor failures
We have lost some scan

Shanghai (Huang Xinyong at Tue Sep 21 00:10:09 METDST 1999) Failure
The cable of the control antenna have broken down.

Effelsberg (A. Kraus at Tue Sep 21 08:14:05 METDST 1999) Success

Torun (K.Borkowski at Tue Sep 21 09:34:23 METDST 1999) Success
No known problems.

Noto (C.Trigilio at Tue Sep 21 12:55:15 METDST 1999) Failure
Noto did not observe because of the failure of the recorder capstain motor

Westerbork (Tony Foley at Tue Sep 21 14:03:54 METDST 1999) Minor failures
Most of the first 2 hours lost because of hardware problems in the array. These were believed fixed about 07:46UT. Also lost from 09:09 to 09:10 with an array restart. 12 telescopes in the array

Jodrell (Alastair Gunn at Wed Sep 29 13:26:30 METDST 1999) Minor failures
Some problems with telescope clock intermittently from about 10:30 UT onwards. Telescope offline between 15:19 and 15:59 UT to solve problem.

Calibration (Jean\-Francois Desmurs at Sun Nov 7 15:44:19 MET 1999) Minor failures
The ANTAB formatted file for this experiment can be found on VLBEER in the sep99 directory under the name: v085k.antab. Read the comments in the file to get further information. (Rem : Jb bad Tsys values.)